OpenStudy (ammarah):

Write an equation of the hyperbola with vertices at (6,-3)(6,1) and foci at (6,-6)(6,4)

OpenStudy (ammarah):

@ganeshie8 so i solved this o ut and for my final answer i got (x-6)^2/16-(y+2)^2/12=1

OpenStudy (ammarah):

the center is (6,-1) a=4, c=2? opens up and down

ganeshie8 (ganeshie8):

it is a vertical hyperbola right ?

ganeshie8 (ganeshie8):

so y^2 term must come first

OpenStudy (ammarah):

ok

OpenStudy (ammarah):



ganeshie8 (ganeshie8):

\(\large \dfrac{(y-(-1))^2}{a^2} - \dfrac{(x-h)^2}{b^2} = 1\)

ganeshie8 (ganeshie8):

\(\large a\) = distance between center and vertex = 1 - -1 = 2

ganeshie8 (ganeshie8):

right ?